A hyperbola with its Center (-3,2), transverse axis parallel to the y-axis, passing through (1,7), the asymptotes are perpendicular to each other.
Answer by lwsshak3(11628)

A hyperbola with its Center (-3,2), transverse axis parallel to the y-axis, passing through (1,7), the asymptotes are perpendicular to each other.
**
Standard form of equation for a hyperbola with vertical transverse axis, as in this case:
(y-k)^2/a^2-(x-h)^2/b^2=1, with (h,k) being the (x,y) coordinates of the center.
..
Took me awhile to figure out that if the asymptotes are perpendicular to each other their slopes must be equal to ±1.
slope=a/b=±1
so, a=b or a=-b
..
Given equation:
(y-2)^2/a^2-(x+3)^2/b^2=1
can be rewritten:
(y-2)^2/a^2-(x+3)^2/a^2=1
using given point on parabola (1,7) to solve for a
(7-2)^2/a^2-(1+3)^2/a^2=1
25/a^2-16/a^2=1
9/a^2=1
a^2=9
a=3
b=a=3
Equation of given hyperbola:
(y-2)^2/9-(x+3)^2/9=1
..
Asymptotes:
Asymptotes are straight lines thru the center of the hyperbola of the standard form:
y=mx+b, m=slope, b=y-intercept
for slope=1
2=1*-3+b
b=5
equation of asymptote
y=x+5
..
for slope=-1
2=-1*-3+b
b=-1
equation of asymptote
y=-x-1
..
See graph as a visual check on answers:
..
y=±(9+(x+3)^2)^.5+2

..